an investor put a total of 3500 in two mutal funds. one fund had a profit of 6% and the other 4% profit, for a total of 170 profit. how much was invested in each fund?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y =
3500
Interest----- 170

Part I 6.00% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 3500 -------------1
Interest on both investments
6.00% x + 4.00% y= 170
Multiply by 2
6 x + 4 y= 17000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -21000.00
Add this to (2)
0 x -2 y= -4000
divide by -2
y = 2000
Part I 6.00% $ 1500
Part II 4.00% $ 2000


CHECK
1500 --------- 6.00% ------- 90.00
2000 ------------- 4.00% ------- 80.00
Total -------------------- 170.00

m.ananth@hotmail.ca